OpenStudy (anonymous):

Find the Taylor Series centered at c=0 for f(x)=40x^(5)/(5+8x) and determine its interval of convergence

OpenStudy (zzr0ck3r):

Do you know the formula for a Maclaurin series?

OpenStudy (anonymous):

First you know that \[ \frac1{1+ u}=\sum_{i=0}^\infty (-1)^n u^n, \quad |u| <1\\ \frac{ 40 x^5}{5+ 8 x}=\frac{ 40 x^5}{5(1+ \frac{8 x}5)}=\frac{ 8 x^5}{(1+ \frac{8 x}5)}\\ \frac{ 1}{1+ \frac{8 x}5}=\sum_{i=0}^\infty (-1)^n\left ( \frac {8x}5\right)^n, \quad |8x/5| <1\\ \] You should be able to finish it now.

OpenStudy (anonymous):

\[ \frac {1} {1 + \frac {8 x} 5} = \sum_ {i = 0}^\infty (-1)^n \frac {8^ n} {5^n} x^n, \quad | x | < 5/8 \\ \frac {8 x^5} {1 + \frac {8 x} 5} = \sum_ {i = 0}^\infty (-1)^n \frac {8^ {n+1}} {5^n} x^{n+5}, \quad | x | < 5/8 \\ \]